OpenStudy (anonymous):

okay im trying to understand 'solving systems by graphing' using linear equations how would i awnser (3,1) ; { x+3y=6 4x-5y=7

OpenStudy (anonymous):

plot these two lines in graph paper and u will see that trhey intersect at the point (3,1)

OpenStudy (anonymous):

i have to tell wether the ordered pair is a soulution to the given system and i dont know how to check that..

OpenStudy (anonymous):

put the order pair in the rigth hand side of both the line and check whether the evaluated value is equal to the RHS or not...

OpenStudy (anonymous):

for the first line 3+3*1=6 RHS for second line 4*3-5*1=12-5=7 =RHS..so it is a solution...

OpenStudy (anonymous):

so but the 3 in the place of x in the x+3y=6 equation

OpenStudy (anonymous):

i think i get it!(:

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

sorry im alittle bad at math.. thaknk you soo much(: